OpenStudy (anonymous):

The difference between a number and eleven is three more than half the number.

OpenStudy (dangerousjesse):

\[x-11=3+.5*x\]

OpenStudy (anonymous):

Thank you! :)

OpenStudy (dangerousjesse):

No problem.

OpenStudy (anonymous):

would that look like x - 11 = x/2 + 3 or x - 11 = x+3/2 + 3?